OpenStudy (anonymous):

help please Solve each equation by the method of completing the square. Show your work. 1. x^2+8x=20 2.x^2-30x=-125

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

@Isaiah.Feynman @ganeshie8 @Hero

OpenStudy (anonymous):

@mathstudent55 @hba

OpenStudy (anonymous):

what do you want

OpenStudy (anonymous):

@ilfy214

OpenStudy (anonymous):

\[x^2+8x=20 \]\[x^2+8x -20 = 0\] Got it from here? \[x^2-30x=-125\]\[x^2-30x+125 = 0\]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Yay! Happy to help :)